OpenStudy (surryyy):

Find the volume of a cylinder whose base has a diameter of 11 and whose height is 12.5. Use π = 3.14 and round your answer to the nearest tenth. Answers: 137.5 392.5 1187.3 953.8

OpenStudy (anonymous):

V= pi*r^2*h V=3.14*(5.5)^2*12.5 V=1187.3

OpenStudy (surryyy):

Thnx so much!